Healthy Foods Inc. sells 50-pound bags of grapes to the military for $10 a bag. The fixed costs of this operation are $80,000, while the variable costs of grapes are $0.10 per pound. 1. a. What is the break-even point in bags? b. Calculate the profit or loss on 12,000 bags and on 25,000 bags.
Answer #1
a) Breakeven point in bags = Fixed cost/ Contribution margin per bag = 80000/(10-50*0.10) = 16000 bags
b) Loss on 12000 bags = Total contribution margin-Fixed costs
[CM per unit = 10-50*0.10 = $5]
Loss on 12000 bags = 12000*5-80000 = $          -20,000
Profit on 25000 bags = 25000*5-80000 = $           45,000